OpenStudy (kabase1234):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . .

OpenStudy (mayankdevnani):

well formula for finding the nth term is :- \[\large \bf a_n=a+(n-1)d\] where a=first terms n=no of that term d=common difference

OpenStudy (kabase1234):

So the first term would be 1?

OpenStudy (kabase1234):

I don't know how to find the commons difference

OpenStudy (mayankdevnani):

common difference is difference between any two consecutive terms

OpenStudy (mayankdevnani):

so here common difference would be -4-1=-5

OpenStudy (mayankdevnani):

got it ?

OpenStudy (kabase1234):

Ohhh okay yes I got it

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

so \[\large \bf a_{12}=1+(12-1)(-5)\]

OpenStudy (mayankdevnani):

find 12th term ?

OpenStudy (kabase1234):

Okay hang on

OpenStudy (kabase1234):

I got -54

OpenStudy (mayankdevnani):

correct

OpenStudy (mayankdevnani):

so in previous question, i told you the formula if we have first term and last term ,we would surely use it now

OpenStudy (mayankdevnani):

remember that formula ?

OpenStudy (kabase1234):

Hang on

OpenStudy (kabase1234):

I can't remember what it was, I'm sorry

OpenStudy (mayankdevnani):

okay, np here is the formula :- \[\large \bf S_n=\frac{n}{2}(a+l)\]

OpenStudy (mayankdevnani):

can you tell me what is the value of n,a and l here ?

OpenStudy (kabase1234):

Is a still 1?

OpenStudy (mayankdevnani):

yep

OpenStudy (kabase1234):

And if we found the 12th term does n change? Or does one of them change to -54?

OpenStudy (kabase1234):

Or is n still 12?

OpenStudy (kabase1234):

Sorry for all the questions in just trying to understand

OpenStudy (mayankdevnani):

yep! n=12 i.e no of terms

OpenStudy (mayankdevnani):

and l=-54 i.e last terms

OpenStudy (mayankdevnani):

*term

OpenStudy (kabase1234):

Okay thank you

OpenStudy (mayankdevnani):

just plug all the values in given equation and solve

OpenStudy (kabase1234):

Okay

OpenStudy (kabase1234):

12

OpenStudy (mayankdevnani):

how ?

OpenStudy (kabase1234):

I'm not sure I think I did it wrong

OpenStudy (mayankdevnani):

haha you did

OpenStudy (kabase1234):

What do I do now?

OpenStudy (mayankdevnani):

\[\large \bf S_{12}=\frac{12}{2}(1-54)\]

OpenStudy (mayankdevnani):

solve it

OpenStudy (kabase1234):

OHHH. I forgot the 54

OpenStudy (kabase1234):

-318

OpenStudy (mayankdevnani):

correct

OpenStudy (kabase1234):

Okay what now? Or did I just finish it??

OpenStudy (mayankdevnani):

haha yes

OpenStudy (kabase1234):

Thank you so much!

OpenStudy (mayankdevnani):

np